# Direct Imitation Learning-based Visual Servoing using the Large Projection Formulation

### v1

@@ 4.3 Merging DL and IL with the large projector
The DL-based backbone detector and the NODE target network are deployed within the robot control loop, as shown in the schematic of our approach (Fig. 3). Given the desired and current image, the perception backbone extracts the visual features, as in (4), which are then used to compute the visual error $\bm{e}$ and its norm $\eta$. These values are needed to compute the primary task in (3). The lower priority (secondary) task considers the corrective velocity $\bm{\sigma}$ as regressed from the NODE target network (9).
In our scheme, the pre-trained YOLO backbone represents the DL detection model. The higher-priority task uses the current features (detected from the camera image) to adapt to changes in the object’s location. This feedback term also ensures convergence to the desired visual features $\bm{f}^{*}$ (i.e., $\bm{e}\rightarrow 0$ for $t\rightarrow+\infty$). At the same time, the NODE network realizes an open-loop IL strategy that lets the robot execute more complex motions without affecting the convergence.
@@ 5.1 Hardware and software components
Our software is implemented using ROS noetic and we use the franka_example_controllers [^1] https://wiki.ros.org/franka_example_controllers package to communicate with the robot. The ROS interface accepts pose-level input, which we obtain by integrating the velocity command, computed as in (2). Furthermore, we experimentally verified that position and orientation tasks are decoupled in (2). This is because the features from the YOLO backbone (obtained from the squared bounding box) do not contain information on the orientation, see Fig. 2. Therefore, as a difference from the general block diagram in Fig. 3, the velocity output of NODE is split into its linear and angular parts; the first actually enters the large projection formulation, whereas the latter goes directly to the robot. This implementation detail implies that the control structure maintains the robot stability in position, while the execution of the orientation is delegated to the mere imitation strategy. Nevertheless, it is worth mentioning that in our experiments, we observe that, with few demonstrations, the robot can perform safe behaviors even in orientation.
We use the realsense2_camera [^2] https://wiki.ros.org/realsense2_camera package to communicate with the camera and run YOLO with the darknet_ros [^3] http://wiki.ros.org/darknet_ros package. We use RGB images with a size of $640\times 480$ pixels and a framerate of $30$ Hz. NODE is implemented and trained in PyTorch. Our open-source code, instructions for execution, and all necessary software dependencies are available at https://github.com/sayantanauddy/d-ilvs.
5.2 YOLO backbone
The YOLO backbone in our setup (see Fig. 3) uses the yolov2-tiny [18] model pretrained on the COCO dataset [52]. The underlying YOLO network can be easily changed to any of the other pre-trained models provided by the darknet_ros3 package. In our experiments, we select the object of interest (e.g., the mouse or the cup) from the list of all objects detected by YOLO and use the detected features of this object for VS.
The visual features (i.e., the corners of the bounding boxes) detected by YOLO are noisy and the features are prone to random shifts even in consecutive frames. Furthermore, the object may not be detected in some frames, leading to missing visual features. To mitigate the effects of these problems, we convert the bounding box detected by YOLO into a square shape with the same center as the original YOLO bounding box and a side length equal to the largest side of the original box. Furthermore, since the object of interest is not subject to rapid movements in the camera frame, we apply average filtering using the past $50$ frames. We use the vertices of the resulting bounding box as features in our VS scheme. In the presentation of our results, such features are denoted with $\text{C}_{i}$ whereas their desired counterparts are $\text{G}_{i}$, with $i=1,\dots,4$.
5.3 Collection of demonstrations
@@
For training NODE, we collect demonstrations via kinesthetic teaching [5]. The object under consideration is placed in a specific location, and a human user physically guides the robot’s end-effector from an initial pose to the desired final pose. We collect two sets of demonstrations corresponding to two different objects. In the first set, a computer mouse is placed on the table with the robot’s camera looking down at the mouse; kinesthetic demonstrations are provided so that the image of the mouse is rotated 90 clockwise in the final pose of the robot. The latter set of demonstrations is collected using a cup. In the initial pose for these demonstrations, the camera looks side-on at a cup on the table; in the final one, the robot’s end-effector is positioned on top of the cup with the camera looking down. See Fig. 4 for a visual reference of the initial and final poses for the mouse and cup demonstrations. Each set contains four demonstrations used to train a NODE (one for each set).
@@ 5.4 NODE training
The NODE target network predicts the derivatives of the inputs, and during training, numerical integration is used to generate trajectories from the predictions [46]. We use a NODE target network with 2 hidden layers containing 256 neurons each and ReLU activations. In each demonstration $n$ and each step $t$ of the training data, the inputs and outputs of the NODE are 10-dimensional, consisting of the upper left and lower right vertices of the bounding box (i.e., the visual features) $\bm{f}^{n}_{t}\in\mathbb{R}^{4}$ in image coordinates normalized to lie within $[0.0,100.0]$, the position (in cm) of the robot’s end-effector in the task space $\bm{p}^{n}_{t}\in\mathbb{R}^{3}$, and the rotation vector $\bm{r}^{n}_{t}\in\mathbb{R}^{3}$ obtained by projecting the orientation quaternions of the end-effector to the local tangent space, as described in Sec. 4.2. Following [46], we scale the rotation vectors by a constant factor of 100.0 so that all input features are of comparable magnitudes. For each recorded demonstration set (mouse and cup), we train a NODE for $2\times 10^{4}$ iterations with a learning rate of $5\times 10^{-4}$ using the loss defined in (8).
@@ 5.5 Evaluation protocol
Our analysis compares the performance of three VS schemes. The first is denoted with Image-based (IIL) and uses a NODE instance trained on the demonstrations to control the robot in an end-to-end fashion. The second one is a classic VS scheme where YOLO provides the required visual feedback; following the literature [34, 35], we call it DVS, as it is a direct approach considering the whole image as input. Finally, our proposed method, augmenting the ILVS scheme [12] with DL-based direct measurement, is called D-ILVS.
We conduct separate experiments for the mouse and the cup. For each experiment, we evaluate the performance of the three schemes for five different object positions: one as in the demonstrations, and four unseen positions, as shown in Fig. 5.
We run each test for $T=700$ time steps, where $T$ is the demonstration length, and measure the norm of the visual error using (6), and the end-effector position and orientation error at the final step of the experiment. The end-effector position error is computed as
@@
Furthermore, for the cup experiment, we measure the success of dropping a small object into the cup at the end of the robot’s motion. Since resetting the robot after every evaluation may introduce some stochasticity, we evaluate each method on each object position three times. Finally, we also perform a qualitative evaluation of our D-ILVS scheme in a cluttered scene where an object is to be dropped into a cup at different positions among several other objects.
@@
(d) Final image with D-ILVS.
@@
6 Experimental results
During the evaluation of the different approaches described in Sec. 5.5, the same trained NODE is used in our D-ILVS approach as well as in IIL where NODE controls the robot in an end-to-end way. This section presents the results of the different approaches in the mouse and cup experiments. Examples of the presented experiments can be viewed at https://fileshare.uibk.ac.at/f/6d36911a884b4636a5cb/.
6.1 Centering the mouse in the image
The goal of the first set of experiments is to move the robot such that the image of the mouse (whose visual features are shown in green in Fig. 6) coincides with its desired image (corresponding to a desired set of features, shown in blue in Fig. 6). Note that the task also requests the robot camera (and end-effector) to rotate in order to reach the desired relative camera-mouse pose. We evaluate each of the three schemes (DVS, IIL, D-ILVS) on five different object positions, as described in Sec. 5.5, and report the overall results in Fig. 7.
Figure 7: Centering the mouse in the image: the average norm of the final visual error (left), position (center), and orientation error (right) achieved with the three schemes, starting from similar trained positions or novel ones. Note that the end-effector position error cannot be computed for the novel positions. Colored boxes show the means and error bars show the 95% confidence interval.
Figure 8: Dropping an object in the cup: the average norm of the final visual error (left), position (center), and orientation error (right) achieved with the three schemes, starting from similar trained positions or novel ones. Note that the end-effector position error cannot be computed for the novel positions. Colored boxes show the means and error bars show the 95% confidence interval.
DVS achieves low visual errors but fails to control the orientation correctly. IIL achieves low orientation errors but cannot adapt to novel object positions leading to a high visual error. In contrast, our D-ILVS approach exhibits a low visual error, can adapt to novel object positions, and controls the robot orientation properly.
More in detail, DVS achieves the lowest visual error as VS with YOLO alone can position the robot camera such that the current and desired visual features match very closely. However, as the features detected by YOLO do not have any information about the real orientation of the mouse, the yaw orientation of the robot’s end-effector does not change at all. As a result, the final orientation reached in this experiment is very different from the desired one (compare the desired pose in Fig. 4 with the final pose achieved by DVS in Fig. 6b). The IIL approach can achieve low orientation error as the underlying NODE has been trained to achieve the demonstrated orientation. However, it cannot compensate for the changes in the novel mouse positions outside the demonstrations, resulting in high visual error. This effect can also be qualitatively observed in Fig. 6c. Our D-ILVS approach can adapt to novel positions of the mouse. At the same time, it utilizes the trained NODE to achieve the correct orientation as shown by the kinesthetic demonstrations. This enables D-ILVS to take advantage of both DVS and IL and achieve low visual and orientation errors, making it the best approach among the ones we evaluate. Fig. 6d shows that D-ILVS achieves a close fit to the desired visual features and the desired orientation.
@@ 6.2 Dropping an object in the cup
Figure 9: Dropping an object in the cup placed in a novel position: robot external view (left column) and the corresponding camera view (right column) as executed by DVS (a,b), IIL (c,d) and D-ILVS (e,f).
The quantitative evaluation is presented in Fig. 8. The different approaches are again evaluated for one trained object position and four novel object positions outside the demonstrations, as described in Sec. 5.5. DVS achieves a low visual error but fails to move the end-effector above the cup and orient it properly, resulting in high position and orientation errors. IIL achieves a low orientation error but cannot adapt to novel object positions leading to a high visual error. As expected, the visual error is much higher for the novel object positions; the orientation error remains low for both trained and novel object positions. D-ILVS, i.e., our approach, achieves low errors for both position and orientation (corresponding to limited visual errors) as it can adapt to novel object positions and orient the robot properly.
A qualitative evaluation is shown in Fig. 9 and confirms the quantitative results. The DVS approach can align the visual features to their desired counterpart (Fig. 9b), but the robot ends up in a completely wrong pose (Fig. 9a). As a result, it cannot drop the grasped object into the cup. IIL attempts to realize the complex trajectory required to execute the dropping task, but it has poor accuracy (Fig. 9c). In particular, for a novel object position, as the one shown in the figure, the robot is unable to adapt and drops the object outside the cup (Fig. 9d). D-ILVS (ours) shows the best performance among all the approaches, as it can cope with changing positions of the cup and drops the object successfully, as shown in Figs. 9e and 9f.
Figure 10 shows the time evolution of the visual error for the trained object position and two novel object positions during a complete experiment (see Fig. 5 for a description of the trained and novel object positions). D-ILVS (ours) achieves a low visual error for both novel as well as trained object positions, whereas IIL achieves much higher visual errors for novel object positions due to its inability to adapt. As expected, the visual error made by DVS stays low throughout.
We also report the success rates of dropping the object into the cup for all methods (see Tab. 1). Each approach is evaluated 15 times as described in Sec. 5.5 (5 object positions for each of the 3 trials per object position) and we report mean values for success. A trial is considered $100\%$ successful if the object drops cleanly into the cup, $50\%$ successful if the object hits the cup’s rim but eventually falls inside, and $0\%$ otherwise. Tab. 1 shows that our D-ILVS approach achieves near-perfect results, while IIL achieves a much lower score since it is unsuccessful in dropping the object into the cup placed at novel positions; DVS is never able to drop the object into the cup and gets a score of 00.
@@
D-ILVS 100.00 95.83 96.67
Figure 11: With D-ILVS, the robot successfully drops the object into the cup placed in different positions of a cluttered table.
@@ 6.3 Handling cluttered scenes
We execute the experiments presented in the previous section in a cluttered setting and qualitatively evaluate the effectiveness of our D-ILVS approach. The cup is placed on the table among several other objects, such as a book, a plate, a clamp, a spatula, and a game controller. In different trials, the location of the cup is varied among the other objects (see Fig. 11).
The pre-trained YOLO object detection model identifies multiple objects in the scene, as shown in Fig. 12. As YOLO provides a list of detected object names and their corresponding visual features, we can easily select the object of interest (the cup, in our case) and use its visual features for VS, see Fig. 12 (right). Once the cup is selected as the desired object, the robot executes the required motion to position its gripper over the cup. Additionally, in a cluttered scene, YOLO offers us the flexibility of easily changing the target object to any other object detected in the scene.
With our D-ILVS approach, the robot successfully drops the object into the cup in a cluttered setting and also adapts its pose to the different locations of the cup, as shown in Fig. 11 (note that the locations of the cup and the corresponding final poses of the robot in these snapshots). Finally, Fig. 13 shows some views of the object being dropped into the cup placed in different cluttered scenes, as seen from the robot’s internal camera.

### v2

@@ 4.3 Merging DL and IL with the large projector
The DL-based detection model and the NODE target network are deployed within the robot control loop, as shown in the schematic of our approach (Fig. 3). Given the desired and current image, the DL-based detector extracts the visual features, as in (4), which are then used to compute the visual error $\bm{e}$ and its norm $\eta$. These values are needed to compute the primary task in (3). The lower priority (secondary) task considers the corrective velocity $\bm{\sigma}$ as regressed from the NODE target network (9).
In our scheme, the pre-trained YOLO model represents the DL-based detector. The higher-priority task uses the current features (detected from the camera image) to adapt to changes in the object’s location. This feedback term also ensures convergence to the desired visual features $\bm{f}^{*}$ (i.e., $\bm{e}\rightarrow 0$ for $t\rightarrow+\infty$). At the same time, the NODE network realizes an open-loop IL strategy that lets the robot execute more complex motions without affecting the convergence.
@@ 5.1 Hardware and software components
Our software is implemented using ROS noetic and we use the franka_example_controllers [^1] https://wiki.ros.org/franka_example_controllers package to communicate with the robot. The ROS interface accepts pose-level input, which we obtain by integrating the velocity command, computed as in (2). Furthermore, we experimentally verified that position and orientation tasks are decoupled in (2). This is because the features from YOLO (obtained from the squared bounding box) do not contain information on the orientation; recall Fig. 2. Therefore, as a difference from the general block diagram in Fig. 3, the velocity output of NODE is split into its linear and angular parts; the first actually enters the large projection formulation, whereas the latter goes directly to the robot. This implementation detail implies that the control structure maintains the robot stability in position, while the execution of the orientation is delegated to the mere imitation strategy. Nevertheless, it is worth mentioning that in our experiments, we observe that, with few demonstrations, the robot can perform safe behaviors even in orientation.
We use the realsense2_camera [^2] https://wiki.ros.org/realsense2_camera package to communicate with the camera and run YOLO with the darknet_ros [^3] http://wiki.ros.org/darknet_ros package, capturing RGB camera images with a size of $640\times 480$ pixels at a framerate of $30$ Hz. We use the standard calibration procedures provided by the cv2444https://opencv.org/ Python library to determine the camera’s intrinsic parameters (consisting of the focal length and the central point). Instead, to determine the extrinsic camera parameters (i.e., the pose of the camera with respect to the robot’s gripper), we use the aruco_ros [^5] https://wiki.ros.org/aruco_ros package and the cv2 library. NODE is implemented and trained in PyTorch. Our open-source code, including the necessary software dependencies and calibration scripts, is available at https://github.com/sayantanauddy/il-dvs.
@@
5.2 YOLO detector
The YOLO detector in our setup (see Fig. 3) uses the yolov2-tiny [18] model pretrained on the COCO dataset [52]. The underlying YOLO network can be easily changed to any of the other pre-trained models provided by the darknet_ros3 package. In our experiments, we select the object of interest (e.g., the mouse or the cup) from the list of all objects detected by YOLO and use the detected features of this object for VS.
The rectangular bounding boxes originally predicted by YOLO often show significant variations in aspect ratio and vertex positions, even in consecutive frames that appear visually identical. These spurious changes cause abrupt jumps in the computed visual error that can disrupt closed-loop control dynamics and impede target convergence. This issue is particularly critical near the end of the robot’s trajectory, where the current bounding box nearly overlaps with the target bounding box, potentially leading to convergence problems. Therefore, we convert the rectangular bounding boxes into squares with side lengths equal to the larger dimension of the original rectangle while maintaining the same center point. To further mitigate noise in the detected visual features (i.e., vertices of the bounding box), we apply average filtering over the past $50$ frames. As the object of interest is not subject to rapid movements in the camera frame, smoothing the error signal facilitates convergence without compromising the robot’s speed, as demonstrated in the supplementary video.
We use the vertices of the resulting bounding box as features in our VS scheme. In the presentation of our results, such features are denoted with “$\text{C}_{i}$” whereas their desired counterparts are “$\text{G}_{i}$”, with $i=1,\dots,4$.
5.3 Collection of demonstrations
Figure 5: Position trajectories of demonstrations provided for the “Centering the mouse in the image” (left) and “Dropping an object in the cup” (right) tasks. Diversity is introduced by starting from different initial poses and also through the differences between each kinesthetic demonstration.
For training NODE, we collect demonstrations via kinesthetic teaching [5]. The object under consideration is placed in a specific location, and a human user physically guides the robot’s end-effector from an initial pose to the desired final pose. We collect two sets of demonstrations corresponding to two different objects. In the first set, a computer mouse is placed on the table with the robot’s camera looking down at the mouse; kinesthetic demonstrations are provided so that the image of the mouse is rotated 90 clockwise in the final pose of the robot. The latter set of demonstrations is collected using a cup. In the initial pose for these demonstrations, the camera looks side-on at a cup on the table; in the final one, the robot’s end-effector is positioned on top of the cup with the camera looking down. See Fig. 4 for a visual reference of the initial and final poses for the mouse and cup demonstrations. Each set contains four demonstrations used to train a NODE (one for each set). Each demonstration consists of a sequence of $500$ steps. In each set, the object’s position remains unchanged but diversity is introduced into the collected data by starting each demonstration from a different pose. Furthermore, manual kinesthetic demonstrations result in different trajectories and also introduce diversity into the training data. The position trajectories of the demonstrations collected for the two tasks are depicted in Fig. 5.
@@ 5.4 NODE training
The NODE target network predicts the derivatives of the inputs, and during training, numerical integration is used to generate trajectories from the predictions [46]. We use a NODE target network with 2 hidden layers containing 256 neurons each and ReLU activations. In each demonstration $n$ and each step $t$ of the training data, the inputs and outputs of the NODE are 10-dimensional, consisting of the upper left and lower right vertices of the bounding box (i.e., the visual features) $\hat{\bm{f}}^{n}_{t}\in\mathbb{R}^{4}$ in image coordinates normalized to lie within $[0.0,100.0]$, the position (in cm) of the robot’s end-effector in the task space $\bm{p}^{n}_{t}\in\mathbb{R}^{3}$, and the rotation vector $\bm{r}^{n}_{t}\in\mathbb{R}^{3}$ obtained by projecting the orientation quaternions of the end-effector to the local tangent space, as described in Sec. 4.2. Following [46], we scale the rotation vectors by a constant factor of 100.0 so that all input features are of comparable magnitudes. For each recorded demonstration set (mouse and cup), we train a NODE for $2\times 10^{4}$ iterations with a learning rate of $5\times 10^{-4}$ using the loss defined in (8).
@@ 5.5 Evaluation protocol
Our analysis compares the performance of three VS schemes. The first is denoted with Image-based (IIL) and uses a NODE instance trained on the demonstrations to control the robot in an end-to-end fashion. The second one is a classic VS scheme where YOLO provides the required visual feedback; following the literature [34, 35], we call it DVS, as it is a direct approach considering the whole image as input. Finally, our proposed method, augmenting the ILVS scheme [12] with DL-based direct measurement, is called IL-DVS. It is worth mentioning that, for a fair comparison, all the approaches are fed with the square and filtered bounding boxes computed as discussed in Sec. 5.2.
We conduct separate experiments for the mouse and the cup. For each experiment, we evaluate the performance of the three schemes for five different object positions: one as in the demonstrations, and four unseen positions, as shown in Fig. 6. We run each test for $T=700$ time steps, where $T$ is the demonstration length, and measure the norm of the visual error. and the end-effector position and orientation error at the final step of the experiment.
Figure 7: Centering the mouse in the image: the average norm of the final visual error (left), position (center), and orientation error (right) achieved with the three schemes, starting from similar trained positions or novel ones. All errors are computed at the final step of the robot’s trajectory. Note that the end-effector position error cannot be computed for the novel positions. Colored boxes show the means and error bars show the 95% confidence interval.
Figure 8: Dropping an object in the cup: the average norm of the final visual error (left), position (center), and orientation error (right) achieved with the three schemes, starting from similar trained positions or novel ones. All errors are computed at the final step of the robot’s trajectory. Note that the end-effector position error cannot be computed for the novel positions. Colored boxes show the means and error bars show the 95% confidence interval.
All errors are computed at the final step of the robot’s trajectory. In particular, the visual error norm $\eta$ is computed using (6). The end-effector position error is computed as
@@
Furthermore, for the cup experiment, we measure the success of dropping a small object into the cup at the end of the robot’s motion. Since resetting the robot after every evaluation may introduce some stochasticity, we evaluate each method on each object position three times. Finally, we also perform a qualitative evaluation of our IL-DVS scheme in a cluttered scene where an object is to be dropped into a cup at different positions among several other objects.
6 Experimental results
During the evaluation of the different approaches described in Sec. 5.5, the same trained NODE is used in our IL-DVS approach as well as in IIL where NODE controls the robot in an end-to-end way. This section presents the results of the different approaches in the mouse and cup experiments. Examples of the presented experiments can be viewed in the attached supplementary video.
6.1 Centering the mouse in the image
@@
(d) Final image with IL-DVS.
@@
The goal of the first set of experiments is to move the robot such that the image of the mouse (whose visual features are shown in green in Fig. 9) coincides with its desired image (corresponding to a desired set of features, shown in blue in Fig. 9). Note that the task also requests the robot camera (and end-effector) to rotate in order to reach the desired relative camera-mouse pose. We evaluate each of the three schemes (DVS, IIL, IL-DVS) on five different object positions, as described in Sec. 5.5, and report the overall results in Fig. 7. DVS achieves low visual errors but fails to control the orientation correctly. IIL achieves low orientation errors but cannot adapt to novel object positions leading to a high visual error. In contrast, our IL-DVS approach exhibits a low visual error, can adapt to novel object positions, and controls the robot orientation properly.
More in detail, DVS achieves the lowest visual error as VS with YOLO alone can position the robot camera such that the current and desired visual features match very closely. However, as the features detected by YOLO do not have any information about the real orientation of the mouse, the yaw orientation of the robot’s end-effector does not change at all. As a result, the final orientation reached in this experiment is very different from the desired one (compare the desired pose in Fig. 4 with the final pose achieved by DVS in Fig. 9b). The IIL approach can achieve low orientation error as the underlying NODE has been trained to achieve the demonstrated orientation. However, it cannot compensate for the changes in the novel mouse positions outside the demonstrations, resulting in high visual error. This effect can also be qualitatively observed in Fig. 9c. Our IL-DVS approach can adapt to novel positions of the mouse. At the same time, it utilizes the trained NODE to achieve the correct orientation as shown by the kinesthetic demonstrations. This enables IL-DVS to take advantage of both DVS and IL and achieve low visual and orientation errors, making it the best approach among the ones we evaluate. Fig. 9d shows that IL-DVS achieves a close fit to the desired visual features and the desired orientation.
@@ 6.2 Dropping an object in the cup
Figure 10: Dropping an object in the cup placed in a novel position: robot external view (left column) and the corresponding camera view (right column) as executed by DVS (a,b), IIL (c,d) and IL-DVS (e,f).
@@
Figure 12: With IL-DVS, the robot successfully drops the object into the cup placed in different positions of a cluttered table.
The quantitative evaluation is presented in Fig. 8. The different approaches are again evaluated for one trained object position and four novel object positions outside the demonstrations, as described in Sec. 5.5. DVS achieves a low visual error but fails to move the end-effector above the cup and orient it properly, resulting in high position and orientation errors. IIL achieves a low orientation error but cannot adapt to novel object positions leading to a high visual error. As expected, the visual error is much higher for the novel object positions; the orientation error remains low for both trained and novel object positions. IL-DVS, i.e., our approach, achieves low errors for both position and orientation (corresponding to limited visual errors) as it can adapt to novel object positions and orient the robot properly.
A qualitative evaluation is shown in Fig. 10 and confirms the quantitative results. The DVS approach can align the visual features to their desired counterpart (Fig. 10b), but the robot ends up in a completely wrong pose (Fig. 10a). As a result, it cannot drop the grasped object into the cup. IIL attempts to realize the complex trajectory required to execute the dropping task, but it has poor accuracy (Fig. 10c). In particular, for a novel object position, as the one shown in the figure, the robot is unable to adapt and drops the object outside the cup (Fig. 10d). IL-DVS (ours) shows the best performance among all the approaches, as it can cope with changing positions of the cup and drops the object successfully, as shown in Figs. 10e and 10f.
Figure 11 shows the time evolution of the visual error for the trained object position and two novel object positions during a complete experiment (see Fig. 6 for a description of the trained and novel object positions). IL-DVS (ours) achieves a low visual error for both novel as well as trained object positions, whereas IIL achieves much higher visual errors for novel object positions due to its inability to adapt. As expected, the visual error made by DVS stays low throughout.
We also report the success rates of dropping the object into the cup for all methods (see Tab. 1). Each approach is evaluated 15 times as described in Sec. 5.5 (5 object positions for each of the 3 trials per object position) and we report mean values for success. A trial is considered $100\%$ successful if the object drops cleanly into the cup, $50\%$ successful if the object hits the cup’s rim but eventually falls inside, and $0\%$ otherwise. Tab. 1 shows that our IL-DVS approach achieves near-perfect results, while IIL achieves a much lower score since it is unsuccessful in dropping the object into the cup placed at novel positions; DVS is never able to drop the object into the cup and gets a score of 00.
@@
IL-DVS 100.00 95.83 96.67
@@ 6.3 Handling cluttered scenes
We execute the experiments presented in the previous section in a cluttered setting and qualitatively evaluate the effectiveness of our IL-DVS approach. The cup is placed on the table among several other objects, such as a book, a plate, a clamp, a spatula, and a game controller. In different trials, the location of the cup is varied among the other objects (see Fig. 12).
The pre-trained YOLO object detection model identifies multiple objects in the scene, as shown in Fig. 13 (left). As YOLO provides a list of detected object names and their corresponding visual features, we can easily select the object of interest (the cup, in our case) and use its visual features for VS, see Fig. 13 (right). Once the cup is selected as the desired object, the robot executes the required motion to position its gripper over the cup. Additionally, in a cluttered scene, YOLO offers us the flexibility of easily changing the target object to any other object detected in the scene.
Figure 14: Two final image frames captured by the robot camera showing that IL-DVS successfully drives the robot to drop the object into the cup in different cluttered environments.
With our IL-DVS approach, the robot successfully drops the object into the cup in a cluttered setting and also adapts its pose to the different locations of the cup, as shown in Fig. 12 (note that the locations of the cup and the corresponding final poses of the robot in these snapshots). Finally, Fig. 14 shows some views of the object being dropped into the cup placed in different cluttered scenes, as seen from the robot’s end-effector camera.